RC meeting and Run Plan, Monday February 6th
From SBS wiki
Return to the RC Meeting Calendar
Contents
RC Report
- https://logbooks.jlab.org/entry/4125474
- https://logbooks.jlab.org/entry/4125522
- RC: Rachel Montgomery, rmontgom@jlab.org
Summary
- Friday swing and owl were productive in production data taking.
- Saturday day and swing were productive in production data taking.
- Sunday owl started having beam issues. Sunday day, swing and Monday owl had almost no beam due to accelerator issues. See RC report for more details.
- Today, Monday, day. Accelerator was shut down at ~8:30am to work on a magnet for Hall D. We went to controlled access ~13:40 to check on:
- GEM cable for plans to insert SBS GEMs in DAQ
- DC power checked Moller magnet - they will post an elog, the issue seems to be with method of ramping up, will post a new procedure
- Target oven heater check.
- Target team were also planning an EPR today.
- This afternoon at around 2pm we got word that Hall D magnet issue was fixed. Ops plan to start testing the beam this afternoon and aim to deliver beam to Halls at some point during swing shift (ETA unkown).
- Ops cancelled RF recovery day tomorrow so please don't expect to access hall tomorrow.
- We started locking up again at 3 and are waiting to do full beam recovery when it is available.
Current Status
- Waiting on beam.
Run Plan
- When ops are ready to send us beam. We will do full beam checkout and then continue production running on polHe3, hopefully 45uA.
Data Collection and Polarisaton Status
- 24.0C of uncorrected charge for this kinematic (26.4C of corrected charge)
Other Topics/Status Updates
- On Saturday we changed ps1=2, checked DAQ set up/event rate (~300-350Hz at 45uA) and live time (~100%) with Alex, any other comments on this?
Shift Schedule
- Shift Schedule Page
- Opening for SL on Fri 10th Feb swing, other than that looking great in coming weeks. Just a handful of spots left to sign up, especially TO - don't miss your chance! Thank you for signing up.
Requested Accesses to the Hall
- Jack would like to collect some cables
Special Requests
Notices for Shift Crew
- Check the How To page
- Pay attention to the beam restoration procedures in the How To
- there is a procedure if beam has been away for 30minutes, and another procedure if machine has been down for tuning or something which may change beam profile. Communicate with MCC to figure out why beam has been down and to figure out which kind of restoration you should do.
- SL:
- production runs (1-1.5hr) and replays (50k replay for each run, ~two runs also with 200k replays per shift if possible)
- TOs:
- NMR every 4 hours
- Note the instructions for performing an NMR if there has been any changes to magnetic fields (https://logbooks.jlab.org/entry/4124321)
- Contact RC any time for any questions